How can i find the vertex of

Y=(x-4)^2 +5

Answer by jim_thompson5910(35256)

The vertex of y = a(x-h)^2+k is (h,k). In the case of y=(x-4)^2+5, we see that h=4 and k=5. So the vertex is (4,5)


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim